OpenStudy (anonymous):

MEDAL AND FAN PLEASE HELP! Determine which system below will produce infinitely many solutions. A) -6x+3y=18 4x-3y=6 B) 2x+4y=24 6x+23y=36 C) 3x-y=14 -9x+3y=-42 D) 5x+2y=13 -x+4y=-6

OpenStudy (anonymous):

@Hotchellerae21 can you help me?

OpenStudy (anonymous):

please :(

OpenStudy (anonymous):

i think B i am not sure

OpenStudy (anonymous):

Okay! Could you help me with one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve the following system of equations: -2x+y=1 -4x+y=-1

OpenStudy (anonymous):

A) (3,1) B) (-1,3) C) (-1,-3) D) (1,3)

OpenStudy (anonymous):

Do you know?

OpenStudy (anonymous):

i think D i don't know if its right or not

OpenStudy (anonymous):

Okay thank you